# CCMNet: Leveraging Calibrated Color Correction Matrices for Cross-Camera Color Constancy

### v1

@@ Abstract
Figure 1: This paper introduces CCMNet, a framework for cross-camera color constancy. CCMNet uses pre-calibrated color correction matrices (CCMs) from camera ISP hardware to train an encoder that generates a camera fingerprint embedding (CFE), capturing the testing camera’s color space. In (A), we show a raw image from a Canon 550D. In (B), we present C5 [6], which generalizes using randomly selected unlabeled images from the test camera—C5’s performance varies depending on the image set. In (C), we show our results, relying only on fixed CCMs in the ISP. Neither method used Canon 550D data during training. Gamma correction was applied for visualization.
@@ 1 Introduction
Computational color constancy ensures that object colors remain consistent under varying lighting conditions [10]. In digital cameras, this is achieved through white balancing, which adjusts raw image colors to simulate neutral lighting [31, 11, 20]. This involves two main steps: illuminant estimation and linear white-balance correction [21].
@@
Illuminant estimation predicts the color of the scene’s light source under the assumption of single-source illumination [31]. The estimated illuminant color is then used in linear white-balance correction to counteract the effects of lighting and camera response biases [26, 17]. These steps are applied early in the image signal processor (ISP) pipeline to raw images [4] and are influenced by the camera’s sensor-specific characteristics, such as response functions and lens properties [52, 2]. These factors complicate the generalization of illuminant estimation algorithms across cameras with varying characteristics [3].
@@
Recent work on illuminant estimation achieves promising results using learning-based models [30, 65, 48, 7]. These models learn a mapping between input image colors and scene illuminant colors using pairs of images and corresponding ground-truth illuminant colors, typically captured by the same camera used in testing [53]. Consequently, most learning-based methods struggle to generalize to new cameras with different characteristics than those used during training [3]. This limitation hinders their practical applicability in manufacturing, as fine-tuning or retraining is necessary for each new camera introduced. Some recent attempts have proposed solutions for improved adaptation through few-shot learning [63], or by using additional unlabeled images captured by the testing camera at inference time to facilitate generalization to the testing camera’s color space [6]. While these methods show promising results, they require capturing new images with the testing camera for adaptation [46], making their performance inherently dependent on the characteristics of those images [6].
@@
While camera ISPs rely on pre-calibrated, camera-specific information to assist in color processing after white balance has been applied, to the best of our knowledge, no prior work has leveraged this calibrated information for the cross-camera color constancy task. Specifically, consumer-grade ISPs rely on calibrated color correction matrices (CCMs) to transform the camera’s raw color space to a device-independent standard color space (e.g., CIE XYZ). These CCMs are carefully calibrated during ISP manufacturing [41], are easily accessible within the ISP’s firmware [33, 21, 17], and are also available in DNG files for post-capture raw rendering [39]. The availability of this information motivated us to utilize this calibrated data to improve cross-camera generalization.
@@
Contribution In this paper, we present CCMNet, a learning-based method for cross-camera color constancy built on the convolutional color constancy (CCC) framework [37, 12, 13, 6]. Our method leverages pre-calibrated color correction matrices (CCMs) available from camera ISPs to transform predefined illuminant colors along the Planckian locus from the device-independent CIE XYZ color space into the raw space of the test camera. These transformed illuminations encode the unique characteristics of the camera’s response function and serve as reference points. The transformed illuminations are compressed into an 8-dimensional embedding, allowing a learnable hypernetwork to adapt to the test camera’s raw color space and generate a camera-specific CCC model tailored to the input image. Additionally, we introduce an augmentation technique that maps training images from a limited set of cameras to imaginary raw spaces, improving generalization. Consequently, CCMNet, which combines a design that dynamically adapts to the raw space of various cameras with a robust data augmentation strategy, accurately estimates illuminant colors for cameras unseen during training (see Fig. 1-C). Our approach is lightweight, accurate, and requires no additional test camera images, unlike prior work [6].
@@ 2 Related Work
A camera’s ISP includes several color processing modules applied in a pipeline fashion. One of the early-stage modules corrects the colors of the captured raw image through two key steps [26, 40, 17]: (1) image white balancing (Sec. 2.1) and (2) transferring the camera raw colors to a standard color space via CCMs (Sec. 2.2).
@@ 2.1 Auto White Balance
As discussed in Sec. 1, auto white balance modules consist of two steps: illuminant estimation and correction. The correction is applied to the linear raw image colors, often using a diagonal correction matrix [10]. Most research focuses on illuminant estimation, which determines the scene’s illuminant color in the camera’s raw space. This can be categorized into learning-free (statistical) methods (e.g., [18, 45, 23, 61, 32, 19, 55, 56, 60]) and learning-based methods (e.g., [12, 36, 48, 59, 7]).
Statistical-based methods rely on specific hypotheses (e.g., gray-world [18], gray-edges [61], etc.) and use statistics derived from the input raw image colors to estimate the illuminant color. As a result, they inherently generalize across different cameras. However, these methods often have limited accuracy and may fail in scenarios where the scene’s illuminant color cannot be reliably inferred from the captured image.
Learning-based methods (e.g., [29, 30, 16, 12, 58, 36, 53, 13, 64, 54, 65, 48, 67, 47]) improve accuracy by training models to map raw colors to illuminant colors. However, most fail to generalize to unseen cameras [3]. Some approaches attempt adaptation via meta-learning and few-shot learning [50], assuming access to a range of illuminant colors from the testing camera [34, 68], or creating generic methods that require fine-tuning on the new camera [14].
Among these efforts, our work falls into a category of methods designed to achieve adaptation without requiring a paired set of images from the test camera, even if that set is small. To this end, the work in [3] (termed SIIE) maps input raw images from different cameras to a learnable working space, reducing disparities in raw color spaces before illuminant estimation. However, this method assumes access to a diverse range of training cameras to effectively learn this mapping, making its accuracy dependent on the variability of the training data. More recently, C5 [6] utilizes additional images captured by the test camera during inference to dynamically generate a CCC model [12, 13]. While this method achieves promising results, its accuracy heavily depends on the characteristics of the additional images provided (see Fig. 1-B).
@@ 2.2 Color Space Transfer via CCMs
Camera sensors exhibit unique spectral sensitivity and bias, resulting in each camera having its own native RGB color space. Camera ISP manufacturers calibrate and apply color correction matrices (CCMs) to facilitate image processing, ensuring an appropriate transformation between the native RGB space and a device-independent standard color space (e.g., CIE XYZ) within the imaging pipeline [15, 40].
Although the transformation between the camera’s raw space and a standard color space is often nonlinear [25, 24, 35, 38], cameras primarily rely on linear transformation matrices due to their simplicity and practical benefits [24]. CCMs are typically calibrated by fitting a $3\times 3$ matrix that maps the raw RGB values of a calibration object (e.g., a color chart) to their corresponding values in a standard color space under an illuminant with a specific correlated color temperature (CCT) [41, 39]. To accommodate diverse lighting conditions, CCMs are precomputed for at least two illuminants (typically for low and high CCTs [51]) and interpolated for intermediate conditions (see Fig. 2).
@@
We leverage this property of CCMs as a bridge to introduce a novel illuminant estimation method that adapts to the color space of unseen cameras. While previous studies have leveraged CCMs for data augmentation [6], none have explored their use during inference to improve illuminant estimation across different cameras. Additionally, we introduce a data augmentation strategy that exploits the linearity of CCMs to enhance generalization. Specifically, we propose a technique to generate imaginary camera images with corresponding CCMs, further improving the robustness and adaptability of our model.
@@ 3 Method
Figure 3: Overview of the CCMNet architecture. (A) Based on CCC [12] and C5 [6], CCMNet includes a network $f$ that generates filters and bias from the $uv$-histograms of the input image. To process query images from diverse camera domains with varying spectral sensitivities, CCMNet uses a camera fingerprint embedding (CFE) as guidance. (B) The CFE for three example cameras (A, B, V)—two real (A, B) and one imaginary (V)—is constructed by mapping predefined illuminants (2500K–7500K along the Planckian locus) from the CIE XYZ space to each camera’s native raw RGB space using calibrated CCMs. These values are converted into a $64\times 64$ histogram and encoded into a 1D vector via a lightweight encoder.
@@ 3.1 Preliminary
As shown in the upper flow of Fig. 3-A, our method is fundamentally based on the CCC framework [12, 13], which transforms the image histogram $N$ into an illuminant heatmap $P$, using a filter $F$ and a bias $B$. CCC reformulates the illumination estimation problem as a coordinate localization task on a log-chroma histogram [22], commonly termed a $uv$-histogram. Specifically, for an image’s RGB pixel $[I_{R},I_{G},I_{B}]$, the log-chroma values, $u$ and $v$, are calculated as follows (pixel coordinate $x$ is omitted for simplicity):
@@
where $F$ and $B$ have the same shape as the $uv$-histogram N, * represents the convolution operation (accelerated using fast Fourier transforms), $\sigma$ represents the softmax operation over the $uv$-coordinate space, and the subscript $i$ denotes the index corresponding to the filter and histogram. Here, $i=0$ refers to the original raw image, while $i\geq 1$ corresponds to augmented images (e.g., texture, edge). We can interpret $P$ as a heatmap of confidence for each $u,v$ coordinate, so the final prediction (^$(\hat{\ell}_{u},\hat{\ell}_{v})$ is expressed as a weighted sum of the coordinates using $P$:
^$\hat{\ell}_{u}=\sum_{u,v}uP(u,v),\;\;\;\hat{\ell}_{v}=\sum_{u,v}vP(u,v).$ (6)
@@
^$\hat{\ell}=\left[\exp\left(-\hat{\ell}_{u}\right),1,\exp\left(-\hat{\ell}_{v}\right)\right],$ (7)
@@
The training objective of CCC is to optimize the filter and bias to minimize the angular error between the predicted illumination RGB and the ground truth illumination in the training dataset. For cross-camera color constancy, C5 [6] proposes a hypernetwork version of CCC that dynamically generates $F$ and $B$ for the test image after analyzing histograms of additional images taken by the same camera.
@@ 3.2 CCMNet
The proposed CCMNet framework is built upon C5 [6]. As mentioned above, C5 is a hypernetwork version of CCC [12, 13], where the network dynamically generates filters and bias. However, unlike C5, CCMNet does not require additional images from the target camera (typically 6–8). Instead, our method leverages two pre-calibrated Color Correction Matrices (CCMs) for low and high correlated color temperatures (CCTs) embedded within the ISP (see Fig. 3-A). These CCMs provide stable guidance via Camera Fingerprint Embedding (CFE) (see Fig. 3-B), ensuring consistent performance across diverse camera domains without extra test images. The core formulation of CCMNet is as follows:
@@ 3.4 Imaginary Camera Augmentation
In prior illuminant estimation research, most data augmentation techniques (e.g., [27, 49, 1, 5]) rely on transferring ground-truth illuminant colors—randomly sampled from a given dataset—to other images within the same dataset (captured by the same camera) using chromatic adaptation. However, this approach is incompatible with our method, which is trained on raw images from different cameras, each with a distinct raw color space. Another augmentation approach [6] leverages camera-specific information and CCMs to perform raw-to-raw augmentation by transferring images from a source camera to a target camera. While promising, this method remains constrained by the limited diversity of training camera raw spaces.
@@
[…]
@@ 4.1 Experimental Setup
The input image and camera-specific raw RGB illuminants (51 colors ranging from 2500K to 7500K, sampled at 100K intervals) are represented as $64\times 64$ $uv$-histograms. The $uv$-ranges are empirically set to [-2.85, 2.85] for the input query image and [-0.5, 1.5] for CFE encoding.
We use the Intel-TAU [44], Gehler-Shi [57], NUS-8 [19], and Cube+ [9] datasets for training and testing. Each dataset includes images captured by distinct cameras, with no overlap between datasets. The number of cameras varies between one (Cube+) and eight (NUS-8).
Following the protocol in C5 [6], we adopt a leave-one-out cross-dataset evaluation approach, where the network is trained on all datasets except the test dataset. For instance, when validating on Gehler-Shi, the network is trained using Intel-TAU, NUS-8, and Cube+, ensuring no camera overlap between training and test datasets. We exclude the Sony-IMX subset of Intel-TAU due to the absence of CCM information, so Intel-TAU is used solely for training.

### v2

@@ Abstract
Figure 1: This paper introduces CCMNet, a framework for cross-camera color constancy. CCMNet uses pre-calibrated color correction matrices (CCMs) from camera ISP hardware to train an encoder that generates a camera fingerprint embedding (CFE), capturing the testing camera’s color space. In (A), we show a raw image from a Canon 550D. In (B), we present C5 [afifi2021cross], which generalizes using randomly selected unlabeled images from the test camera—C5’s performance varies depending on the image set. In (C), we show our results, relying only on fixed CCMs in the ISP. Neither method used Canon 550D data during training. Gamma correction was applied for visualization.
@@ 1 Introduction
Computational color constancy ensures that object colors remain consistent under varying lighting conditions [barnard1995computational]. In digital cameras, this is achieved through white balancing, which adjusts raw image colors to simulate neutral lighting [gijsenij2011computational, barnard2002comparison, cheng2015beyond]. This involves two main steps: illuminant estimation and linear white-balance correction [delbracio2021mobile].
@@
Illuminant estimation predicts the color of the scene’s light source under the assumption of single-source illumination [gijsenij2011computational]. The estimated illuminant color is then used in linear white-balance correction to counteract the effects of lighting and camera response biases [florin2009color, brown2023color]. These steps are applied early in the image signal processor (ISP) pipeline to raw images [afifi2019color] and are influenced by the camera’s sensor-specific characteristics, such as response functions and lens properties [nguyen2014raw, afifi2021semi]. These factors complicate the generalization of illuminant estimation algorithms across cameras with varying characteristics [afifi2019sensor].
@@
Recent work on illuminant estimation achieves promising results using learning-based models [gijsenij2010generalized, yu2020cascading, lo2021clcc, afifi2025optimizing]. These models learn a mapping between input image colors and scene illuminant colors using pairs of images and corresponding ground-truth illuminant colors, typically captured by the same camera used in testing [oh2017approaching]. Consequently, most learning-based methods struggle to generalize to new cameras with different characteristics than those used during training [afifi2019sensor]. This limitation hinders their practical applicability in manufacturing, as fine-tuning or retraining is necessary for each new camera introduced. Some recent attempts have proposed solutions for improved adaptation through few-shot learning [xiao2020multi], or by using additional unlabeled images captured by the testing camera at inference time to facilitate generalization to the testing camera’s color space [afifi2021cross]. While these methods show promising results, they require capturing new images with the testing camera for adaptation [li2023ranking], making their performance inherently dependent on the characteristics of those images [afifi2021cross].
@@
While camera ISPs rely on pre-calibrated, camera-specific information to assist in color processing after white balance has been applied, to the best of our knowledge, no prior work has leveraged this calibrated information for the cross-camera color constancy task. Specifically, consumer-grade ISPs rely on calibrated color correction matrices (CCMs) to transform the camera’s raw color space to a device-independent standard color space (e.g., CIE XYZ). These CCMs are carefully calibrated during ISP manufacturing [karaimer2018improving], are easily accessible within the ISP’s firmware [hasinoff2016burst, delbracio2021mobile, brown2023color], and are also available in DNG files for post-capture raw rendering [adobe2023digital]. The availability of this information motivated us to utilize this calibrated data to improve cross-camera generalization.
@@
Contribution In this paper, we present CCMNet, a learning-based method for cross-camera color constancy built on the convolutional color constancy (CCC) framework [hubel2007white, barron2015convolutional, barron2017fast, afifi2021cross]. Our method leverages pre-calibrated color correction matrices (CCMs) available from camera ISPs to transform predefined illuminant colors along the Planckian locus from the device-independent CIE XYZ color space into the raw space of the test camera. These transformed illuminations encode the unique characteristics of the camera’s response function and serve as reference points. The transformed illuminations are compressed into an 8-dimensional embedding, allowing a learnable hypernetwork to adapt to the test camera’s raw color space and generate a camera-specific CCC model tailored to the input image. Additionally, we introduce an augmentation technique that maps training images from a limited set of cameras to imaginary raw spaces, improving generalization. Consequently, CCMNet, which combines a design that dynamically adapts to the raw space of various cameras with a robust data augmentation strategy, accurately estimates illuminant colors for cameras unseen during training (see Fig. 1-C). Our approach is lightweight, accurate, and requires no additional test camera images, unlike prior work [afifi2021cross].
@@ 2 Related Work
A camera’s ISP includes several color processing modules applied in a pipeline fashion. One of the early-stage modules corrects the colors of the captured raw image through two key steps [florin2009color, karaimer2016software, brown2023color]: (1) image white balancing (Sec. 2.1) and (2) transferring the camera raw colors to a standard color space via CCMs (Sec. 2.2).
@@ 2.1 Auto White Balance
As discussed in Sec. 1, auto white balance modules consist of two steps: illuminant estimation and correction. The correction is applied to the linear raw image colors, often using a diagonal correction matrix [barnard1995computational]. Most research focuses on illuminant estimation, which determines the scene’s illuminant color in the camera’s raw space. This can be categorized into learning-free (statistical) methods (e.g., [buchsbaum1980spatial, land1977retinex, finlayson2004shades, van2007edge, gijsenij2011improving, cheng2014illuminant, qian2018revisiting, qian2019finding, ulucan2024multi]) and learning-based methods (e.g., [barron2015convolutional, hu2017fc4, lo2021clcc, tang2022transfer, afifi2025optimizing]).
Statistical-based methods rely on specific hypotheses (e.g., gray-world [buchsbaum1980spatial], gray-edges [van2007edge], etc.) and use statistics derived from the input raw image colors to estimate the illuminant color. As a result, they inherently generalize across different cameras. However, these methods often have limited accuracy and may fail in scenarios where the scene’s illuminant color cannot be reliably inferred from the captured image.
Learning-based methods (e.g., [gehler2008bayesian, gijsenij2010generalized, bianco2015color, barron2015convolutional, shi2016deep, hu2017fc4, oh2017approaching, barron2017fast, xu2020end, qian2017recurrent, yu2020cascading, lo2021clcc, yue2023color, kim2021large, kim2024attentive, li2024nightcc]) improve accuracy by training models to map raw colors to illuminant colors. However, most fail to generalize to unseen cameras [afifi2019sensor]. Some approaches attempt adaptation via meta-learning and few-shot learning [mcdonagh2018formulating], assuming access to a range of illuminant colors from the testing camera [hernandez2020multi, yue2024effective], or creating generic methods that require fine-tuning on the new camera [bianco2019quasi].
Among these efforts, our work falls into a category of methods designed to achieve adaptation without requiring a paired set of images from the test camera, even if that set is small. To this end, the work in [afifi2019sensor] (termed SIIE) maps input raw images from different cameras to a learnable working space, reducing disparities in raw color spaces before illuminant estimation. However, this method assumes access to a diverse range of training cameras to effectively learn this mapping, making its accuracy dependent on the variability of the training data. More recently, C5 [afifi2021cross] utilizes additional images captured by the test camera during inference to dynamically generate a CCC model [barron2015convolutional, barron2017fast]. While this method achieves promising results, its accuracy heavily depends on the characteristics of the additional images provided (see Fig. 1-B).
@@ 2.2 Color Space Transfer via CCMs
Camera sensors exhibit unique spectral sensitivity and bias, resulting in each camera having its own native RGB color space. Camera ISP manufacturers calibrate and apply color correction matrices (CCMs) to facilitate image processing, ensuring an appropriate transformation between the native RGB space and a device-independent standard color space (e.g., CIE XYZ) within the imaging pipeline [bianco2013color, karaimer2016software].
Although the transformation between the camera’s raw space and a standard color space is often nonlinear [finlayson2015color, finlayson2020designing, hong2001study, hung1993colorimetric], cameras primarily rely on linear transformation matrices due to their simplicity and practical benefits [finlayson2020designing]. CCMs are typically calibrated by fitting a $3\times 3$ matrix that maps the raw RGB values of a calibration object (e.g., a color chart) to their corresponding values in a standard color space under an illuminant with a specific correlated color temperature (CCT) [karaimer2018improving, adobe2023digital]. To accommodate diverse lighting conditions, CCMs are precomputed for at least two illuminants (typically for low and high CCTs [mcelvain2013camera]) and interpolated for intermediate conditions (see Fig. 2).
@@
We leverage this property of CCMs as a bridge to introduce a novel illuminant estimation method that adapts to the color space of unseen cameras. While previous studies have leveraged CCMs for data augmentation [afifi2021cross], none have explored their use during inference to improve illuminant estimation across different cameras. Additionally, we introduce a data augmentation strategy that exploits the linearity of CCMs to enhance generalization. Specifically, we propose a technique to generate imaginary camera images with corresponding CCMs, further improving the robustness and adaptability of our model.
@@ 3 Method
Figure 3: Overview of the CCMNet architecture. (A) Based on CCC [barron2015convolutional] and C5 [afifi2021cross], CCMNet includes a network $f$ that generates filters and bias from the uvuv-histograms of the input image. To process query images from diverse camera domains with varying spectral sensitivities, CCMNet uses a camera fingerprint embedding (CFE) as guidance. (B) The CFE for three example cameras (A, B, V)—two real (A, B) and one imaginary (V)—is constructed by mapping predefined illuminants (2500K–7500K along the Planckian locus) from the CIE XYZ space to each camera’s native raw RGB space using calibrated CCMs. These values are converted into a $64\times 64$ histogram and encoded into a 1D vector via a lightweight encoder.
@@ 3.1 Preliminary
As shown in the upper flow of Fig. 3-A, our method is fundamentally based on the CCC framework [barron2015convolutional, barron2017fast], which transforms the image histogram $N$ into an illuminant heatmap $P$, using a filter $F$ and a bias $B$. CCC reformulates the illumination estimation problem as a coordinate localization task on a log-chroma histogram [finlayson2001color], commonly termed a $uv$-histogram. Specifically, for an image’s RGB pixel $I_{R},I_{G},I_{B}]$, the log-chroma values, $u$ and $v$, are calculated as follows (pixel coordinate $x$ is omitted for simplicity):
@@
where $F$ and $B$ have the same shape as the $uv$-histogram N, * represents the convolution operation (accelerated using fast Fourier transforms), $\sigma$ represents the softmax operation over the $uv$-coordinate space, and the subscript $i$ denotes the index corresponding to the filter and histogram. Here, $i=0$ refers to the original raw image, while $i\geq 1$ corresponds to augmented images (e.g., texture, edge). We can interpret $P$ as a heatmap of confidence for each $u,v$ coordinate, so the final prediction $(\hat{\ell}_{u},\hat{\ell}_{v})$ is expressed as a weighted sum of the coordinates using $P$:
$\hat{\ell}_{u}=\sum_{u,v}uP(u,v),\;\;\;\hat{\ell}_{v}=\sum_{u,v}vP(u,v).$ (6)
@@
$\hat{\ell}=\left[\exp\left(-\hat{\ell}_{u}\right),1,\exp\left(-\hat{\ell}_{v}\right)\right],$ (7)
@@
The training objective of CCC is to optimize the filter and bias to minimize the angular error between the predicted illumination RGB and the ground truth illumination in the training dataset. For cross-camera color constancy, C5 [afifi2021cross] proposes a hypernetwork version of CCC that dynamically generates $F$ and $B$ for the test image after analyzing histograms of additional images taken by the same camera.
@@ 3.2 CCMNet
The proposed CCMNet framework is built upon C5 [afifi2021cross]. As mentioned above, C5 is a hypernetwork version of CCC [barron2015convolutional, barron2017fast], where the network dynamically generates filters and bias. However, unlike C5, CCMNet does not require additional images from the target camera (typically 6–8). Instead, our method leverages two pre-calibrated Color Correction Matrices (CCMs) for low and high correlated color temperatures (CCTs) embedded within the ISP (see Fig. 3-A). These CCMs provide stable guidance via Camera Fingerprint Embedding (CFE) (see Fig. 3-B), ensuring consistent performance across diverse camera domains without extra test images. The core formulation of CCMNet is as follows:
@@ 3.4 Imaginary Camera Augmentation
In prior illuminant estimation research, most data augmentation techniques (e.g., [fourure2016mixed, lou2015color, abdelhamed2021leveraging, Afifi2020CIEXN]) rely on transferring ground-truth illuminant colors—randomly sampled from a given dataset—to other images within the same dataset (captured by the same camera) using chromatic adaptation. However, this approach is incompatible with our method, which is trained on raw images from different cameras, each with a distinct raw color space. Another augmentation approach [afifi2021cross] leverages camera-specific information and CCMs to perform raw-to-raw augmentation by transferring images from a source camera to a target camera. While promising, this method remains constrained by the limited diversity of training camera raw spaces.
@@
$I_{c}^{V}$ $=\int S(\lambda)R(\lambda)(\alpha Q_{c}^{A}+(1-\alpha)Q_{c}^{B})(\lambda)\,d\lambda$ (12)
@@ 4.1 Experimental Setup
The input image and camera-specific raw RGB illuminants (51 colors ranging from 2500K to 7500K, sampled at 100K intervals) are represented as $64\times 64$ uvuv-histograms. The $uv$-ranges are empirically set to [-2.85, 2.85] for the input query image and [-0.5, 1.5] for CFE encoding.
We use the Intel-TAU [laakom2021intel], Gehler-Shi [shi2000re], NUS-8 [cheng2014illuminant], and Cube+ [banic2017unsupervised] datasets for training and testing. Each dataset includes images captured by distinct cameras, with no overlap between datasets. The number of cameras varies between one (Cube+) and eight (NUS-8).
Following the protocol in C5 [afifi2021cross], we adopt a leave-one-out cross-dataset evaluation approach, where the network is trained on all datasets except the test dataset. For instance, when validating on Gehler-Shi, the network is trained using Intel-TAU, NUS-8, and Cube+, ensuring no camera overlap between training and test datasets. We exclude the Sony-IMX subset of Intel-TAU due to the absence of CCM information, so Intel-TAU is used solely for training.